Fermi National Laboratory


Accelerator Update

Friday June 4
The day shift began with Operations monitoring store 3554 and the Recycler stash, with the Antiproton Source (Pbar) in access for repairs, and with MiniBooNE taking beam.

A Pbar technician reported that he had finished the LCW work on the Pbar Accumulator quadrupole (A:LQ) at 8:35 AM.

After operators resolved some problems with the Pbar voltage bus, they reestablished stacking at 9:48 AM.

Operations terminated store 3554 at 11:04 AM.

Operators prepared the Collider Detector at Fermilab (CDF) and D-zero (D0) for access at 11:17 AM.

Operations halted stacking at 11:25 AM while Pbar experts conducted a thirty-five minute lattice ramping study.

D0 and CDF came out of access around 1 PM.

At 2:11 PM, a Booster kicker (MKS02) tripped off and wouldn't reset. Operators informed a Booster expert who had the kicker back on line by 3:05 PM. Stacking and beam to MiniBooNE resumed at 3:08 PM.

Operations loaded the Tevatron (TeV) with a 12x0 study-store at 7:40 PM.

TeV studies ended at 8:39 PM and then Operations put the TeV through a wet squeeze.

Operations put the TeV into shot setup at 9:18 PM.

Operations established store 3558 at 11:53 PM with an initial luminosity of 55.62E30.

Saturday June 5
The midnight shift began with Operations monitoring store 3558 and the stash, with Pbar stacking, and with MiniBooNE taking beam.

Operations began setting up to send beam to the Meson MCenter experiment at 1:08 AM.

Operations started a rapid stash of antiprotons from Pbar to the Recycler at 3:29 AM. This halted beam to Switchyard and the Meson experiments.

At 7 AM, CDF experimenters called complaining about colliding beam problems. Operators adjusted collimators and had the problem resolved by 7:30 AM.

At 8:41 AM, CDF called again about colliding beam problems. Operations adjusted the TeV dampers.

At 10:35 AM, operators noticed problems with the colliding beam and tried tuning, but had no success. CDF turned off their silicon detector. Operators called in expert help at 10:58 AM.

At 12:43 PM, operators again had to tune to fix TeV colliding beam problems.

At 2:57 PM, the TeV store suffered from an abort. (After eight plus hours of investigation, neither the TeV experts nor Operations found a cause for the abort.)

At 10:17 PM, three Booster RF stations (BRF 9, 11, & 14) tripped off. Stations 9 and 14 would not reset. Operators moved the power sources for certain devices and got the stations to reset.

Pbar resumed stacking at 11:03 PM.

Operations put the TeV into shot setup at 11:23 PM. This halted the Switchyard beam.

Sunday June 6
The midnight shift began with the TeV in shot setup, with Operations monitoring the stash, and with MiniBooNE taking beam.

Operations established store 3560 at 2:12 AM with an initial luminosity of 73.02E30.

Operations reported at 2:57 AM that they were trying to tune away colliding beam problems. CDF turned on their detector at 3:22 AM.

Operations began stashing antiprotons to the Recycler at 4:28 AM. The transfers stopped and stacking resumed at 5:11 AM.

Operators adjusted a TeV collimator to keep the colliding beam in line.

At 11:41 AM, there was a spike in the TeV beam that caused some CDF and D0 devices to trip off. Operations investigated.

Monday June 7
The midnight shift began with Operations monitoring the store and stash, with a Pbar expert getting ready to begin a study period, and with MiniBooNE and MCenter taking beam.

Pbar completed their studies and resumed stacking at 4:22 AM.

The MCenter experiment finished using beam at 7:12 AM.

At 7:36 AM, operators made a TeV collimator adjustment to fix a colliding beam problem.

Plans for Today
Operations will terminate the present store at noon. A TeV expert will conduct a Pot study before the next store gets established. The TeV, Main Injector, Pbar, and the Recycler have been and will continue to conduct studies with the idea of having a mixed mode transfer of antiprotons from both Pbar and Recycler into the TeV later this week.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 6/7/2004   email Fermilab